OpenStudy (anonymous):

A physician saw 18 cases of chicken pox in the first week of an outbreak. Each week, he saw 6 more cases than the week before. After 10 weeks, how many cases of chicken pox did the physician see altogether? Write a sequence that shows the number of cases seen for each of the 10 weeks during the outbreak.

OpenStudy (anonymous):

18 + (0)6, 18 + (1)6, 18 + (2)6, 18 + (3)6, 18 + (4)6, 18 + (5)6, 18 + (6)6, 18 + (7)6, 18 + (8)6, 18 + (9)6 10th week is 18 + (9)6 = 18 + 54 = 72

OpenStudy (anonymous):

All together is adding all the weeks: (18)(10) + (6)(9)(10)/2

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @Owa_Chan

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

The last part of that answer is 6 times the sum of 1 thru 9: sum of 1 thru n is n(n + 1)/2

OpenStudy (anonymous):

So, the sum of 1 thru 9 is: 9(9 + 1)/2